1) Consider a 10-year bond trading at $1150 today. The bond has a face value of $1,000, and has a coupon rate of 8%. Coupons are paid semiannually, and the next coupon payment is exactly 6 months from now. What is the bond's yield to maturity?

2)A coupon-paying bond is trading below par. How does the bond's YTM compare to its coupon rate?

a. Need more info

b. YTM = Coupon Rate

c. YTM > Coupon Rate

d. YTM < Coupon Rate

3)A coupon-paying bond is trading at par. How does the bond's current yield compare to its coupon rate?

a. Need more info

b. YTM = Coupon Rate

c. YTM > Coupon Rate

d. YTM < Coupon Rate

4) Consider a bond with a coupon rate of 6% and a face value of $1,000. Coupons are paid semi-annually. Suppose there are 94 days to the next coupon payment date. Assuming a 30/360 day-count convention, what is the accrued interest on this bond today?

5) Consider a bond with a coupon rate of 7% and a face value of $1,000. Coupons are paid semi-annually. Suppose there are 52 days to the next coupon payment date, beyond which there are 2 years left to maturity (so that there are in total 1+2*2 number of coupon payments left). The bond is currently trading at a YTM of 4%. Assuming a 30/360 day-count convention, what is the bond's full (dirty) price?

6) When buying a bond, an investor pays the clean (or the quoted) price.

a.True

b. False

7)Coupon-bearing Treasury securities pay coupons semi-annually.

a. True

b.False

8)

Calculate the full (dirty) price of a bond, assuming the following.

Transaction settlement date is July 11th, 2018.

Bond maturity date is September 30th, 2020.

Coupon rate is 5%, and coupons are paid semi-annually.

The bond is trading at a YTM of 8%.

The day-count convention is 30/360.

Redemption value is 100.

Round your answer to the nearest cent (2 decimal places).

9)

Consider a corporate bond with a face value of $1,000, 2 years to maturity and a coupon rate of 6%. Coupons are paid semi-annually. The next coupon payment is to be made exactly 6 months from today. What is this bond's price assuming the following spot rate curve.

6-month spot rate: 3%.

12-month: 5%.

18-month: 5.5%.

24-month: 5.8%.

Round your answer to the nearest cent (2 decimal places).

10)

Consider a corporate bond with a face value of $1,000, 2 years to maturity and a coupon rate of 5%. Coupons are paid semi-annually. The next coupon payment is to be made exactly 6 months from today. What is this bond's YTM assuming the following spot rate curve.

6-month spot rate: 4%.

12-month: 5%.

18-month: 5.5%.

24-month: 5.8%.

11) What is the current yield of a 6-year Treasury note with a coupon rate of 5%, a face value of $100, and is currently trading at 100:10?
Answer #1

Answer to Question No. 1

Yeild to maturity (YTM) can be calculated using two methods:

a) Approximation Method

b) IRR Method

Both the methods gives approximately same results.

we are using approximation method to solve this question.

YTM= (C+(F-P)/n) / ((F+P)/2)

where C= Coupon Payment/Interest Payment

F= Face Value

P=Price

n=years to maturity

Putting Values in the equation above:

YTM= (1000*4%+(1000-1150)/20) / ((1000+1150)/2)

Coupon rate is 8% for an year; so 4% semi annually

years to maturity is 10 years; 20 semi annual

YTM=(40+(-7.5))/1075

YTM=3.02% semi annually

So YTM= 3.02%*2= 6.05% annually

Answer to question no.2

If a coupon paying bond is trading below par ; lets say 999

putting the above price in above formulla

YTM= (1000*4%+(1000-999)/20)/ ((1000+999)/2)

YTM= 4.01% semi annually

YTM= 8.01% annually

So correct answer is (C) YTM> Coupon Rate

Answer to Question No. 3

If a bond is trading at par; i.e 1000

YTM=(1000*4%+(1000-1000)/20)/((1000+1000)/2)

YTM=4 % semi annually

YTM= 8% annually

SO correct part is (b) YTM=Coupon Rate

Answer to Question 4

coupon rate= 6% annually , so 3% semi annually

Assumed 360 days in an year, so 180 days in 6 months

94 days left for the next coupon payment, it means we have already earned interest for 86 days (180 days-94 days)

Accrued Interest= 1000*3%*86/180

Accried Intererst = $14.33